### v1

# On the first-order theories of quaternions and octonions

Abstract.
Let $L$ be the language of rings. We provide an axiomatization of the $L$-theories of quaternions and octonions and we characterize the models of mentioned theories: they coincide, up to isomorphism, to quaternion and octonion algebras over a real closed field, respectively. We prove these theories are complete, model complete and they do not have quantifier elimination. Then, we focus on the class of ordered polynomials. Over $\mathbb{H}$ and $\mathbb{O}$ these polynomials are of special interest in hypercomplex analysis slice they are slice regular. We deduce some fundamental properties of the zero locus of ordered polynomials from completeness and we prove the failure of qantifier elimination for the fragment of ordered formulas as well.
Key words and phrases:
Centrally finite alternative division rings; algebraic closure; completeness; model completeness; fundamental theorem of algebra; polynomials of hypercomplex variables; quaternions; octonions; real closed fields
2010 Mathematics Subject Classification:
03C10, 03C98, 16K20, 17A35, 30G35, 14P10
The author is supported by GNSAGA of INDAM
Introduction
In real and complex geometry, both algebraic and analytic, model theory constitutes a fundamental point of view on the subject. In the algebraic setting quantifier elimination established by Chevalley in the complex case [Che43] and by Tarski [Tar51] in the real case constituted mild stones for a further development of algebraic geometry in these fields. In particular, modern real algebraic and semialgebraic geometry was born essentially after Tarski’s result whose consequences are nowadays very clear in algoritmic algebraic geometry as well. For a very complete treatment on these topics for the real algebraic case we refer to [BCR98, BPR06].
In the real analytic setting an example by Osgood [Osg16] shows the existence of an analytic map whose image is not semianalytic, which implies the nonexistence of a symilar result as Tarski’s elimination in the real analytic setting. More in detail, a class of sets stable by boolean operations and by projection needs extra structure to be defined and is called subanalytic. After its definition by A. Gabrielov, the class of subanalytic sets and functions has been deeply studied from the point of view of geometry, see [Hir73, BM88]. From the 80s model theoriests deeply focused on ordered structures, in particular on those sharing many finiteness properties with the ordered field of the reals which are called o-minimal structures. After their definition by Pillay and Steinhord [PS86], the interactions of these structures with real analytic geometry has been deeply exploited by van den Dries and its collaborators, see [DM96, Dri98]. In particular, subanalytic sets are a special case of o-minimal structure. We recall that recently new applications of model theory and o-minimality appeared in complex analytic geometry as well, more precisely in Hodge theory, see the brand new papers [BKT20, BBT23, BKU24].
The aim of this paper is to introduce and develop basic model theoretical properties of the division rings of quaternions and octonions with applications in hypercomplex analysis and geometry. This quite recent subject studies the analog of complex analysis in the more general setting of real alternative ***-algebras of finite dimension. After many attempts during the 20th century, for instance by Fueter, in 2007 Gentili and Struppa [GS07] first defined a notion of regular function $f:\mathbb{H}\to\mathbb{H}$, they called slice regular function, that generalizes to quaternions the classical concept of complex holomorphic function including, for the first time, quaternionic polynomials with coefficients on one side. This new notion of quaternionic regularity has generated a great deal of interest and led both to a deep development of the theory in the quaternionic case, an almost complete reference is [GSS13], and to further extensions over very general real alternative ***-algebras of finite dimension [GP11], including octonions and Clifford algebras. Then, Ghiloni and Perotti [GP22] developed the theory of slice regular functions in several variables. A remarkable fact is slice regular functions are real analytic functions, with respect to real coordinates, and polynomials with ordered variables and coefficients on one side are slice regular. Consequently, a very recent topic to investigate is the class of algebraic subsets of $\mathbb{H}^{n}$ and $\mathbb{O}^{n}$ defined by slice regular polynomial equations. In particular, first results in this direction concern the description of the zero locus of a slice regular polynomial, see [GS08, GP22], and the approach coming from commutative algebra, for example the quaternionic versions of Hilbert’s Nullstellensatz in [GSV24].
These latter subjects strongly motivate the study of elimination results as Chevalley’s and Tarski’s theorem in the quaternionic and octonionic setting. In our paper we provide an axiomatization of the theories of quaternions and octonions in Definitions 1.4&1.5, we prove these theories are complete and model complete in Theorems 2.4&2.5 but they do not afmit quantifier elimination with respect to the language of rings in Corollary 2.12. We point out that, up to author’s knowledge, such a model theory approach on the subject is appearing here for the first time. The main point of our approach concern the possibility of expressing real coordinates in a first-order way, thus results from real algebraic geometry are available in the quaternionic and octonionic theories. Then, we introduce sundamental properties of what we call ordered polynomials with coefficients in every model of mentioned first-order theories. We point out that ordered polynomials over $\mathbb{H}$ and $\mathbb{O}$ conicides with mentioned slice regular polynomials.
This paper proposes a new perspective in the study of slice regular functions as well. As already mentioned, slice regular functions are in particular real analytic functions with respect to real coordinates. Thus, the class of definable slice regular functions actually coincides with the class of slice regular functions which are Nash with respect to real coordinates. In particular, the last equivalence and the definition of slice regular function suggest the possibility of exploiting methods from o-minimality as well in the subject. For this purpose, the results of Peterzil and Starchenko [PS01, PS03] on definable complex analysis using o-minimality are fundamental starting points to investigate the quaternionic and octonionic cases for slice regular functions.
1. Axiomatizations of ACQ &\&& ACO
Let $L:=\{+,-,\cdot,0,1\}$ denote the language of rings. The aim of this section is to provide two classes of first-order $L$-structures one containing $\mathbb{H}$, the algebra of quaternions, and the other containing $\mathbb{O}$, the algebra of octonions. At the end of the day, those structures will exactly coincide with the models of the $L$-theory of quaternions $\textnormal{Th}_{L}(\mathbb{H})$ and of the $L$-theory of octonions $\textnormal{Th}_{L}(\mathbb{O})$, respectively.
Let $D$ be a ring with unity $1\neq 0$. We say that $D$ is alternative if for every $a,b\in D$ the following holds: $a(ab)=aab$ and $(ab)b=a(bb)$. Hence, if $D$ is alternative no parenthesis are needed in the expressions $a^{n}b$ and $ab^{n}$. We say that $D$ is a division ring if for every $a,b\in D$, with $a\neq 0$, there are unique $c,d\in D$ such that $ac=b$ and $da=b$. In particular, if $D$ is a division ring, then every non-null element $a\in D$ admits unique two-sided inverses, that is, there are unique $c,d\in D$ such that $ac=1$ and $da=1$. Denote by
$R:=\{a\in D\,|\,(ab)c=a(bc)=(ca)b,\,\forall b,c\in D\}$
the center of $D$, that is, the subring of $D$ constituting of all elements both associating and commuting with any other element of $D$.
Remark 1.1.
If $D$ is an division ring, then its center $R$ is a field. Indeed, by definition $R$ is a commutative and associative ring. We are left to prove that for every $a\in R$, with $a\neq 0$, the two-sided inverses, namely the unique $c,d\in D$ such that $ac=1$ and $da=1$, actually coincide and are contained in $R$. By associativity and commutativity of $a\in R$ with respect to any element of $D$ we obtain that $d=d(ac)=(da)c=c$, thus the two-sided inverses $c,d\in D$ of $a$ actually coincide. In addition, for every $b\in D$ we have $cb=(cb)(ac)=((cb)a)c=(a(cb))c=((ac)b)c=bc$ and for every $e,f\in D$ we have $(ce)f=c((e(ac))f)=c(((ea)c)f)=c(((ae)c)f)=(ca)((ec)f)=((ec)f)=(ce)f$, as desired.
If $D$ is a division ring, by Remark 1.1, $D$ is in particular a vector space over its center $R$. We say that $D$ is centrally finite if its dimension as a $R$-vector space is finite.
Example 1.2.
Let $F$ be a field of characteristic $\neq 2$. Denote by $\mathbb{H}_{F}$ and $\mathbb{O}_{F}$ the rings of quaternions and octonions over $F$, respectively. We refer to [Ebb+91, Sch66] for a precise definition of those rings. If $F$ is an ordered field, then $\mathbb{H}_{F}$ and $\mathbb{O}_{F}$ are centrally finite alternative division rings. In particular, $\mathbb{H}_{F}$ and $\mathbb{O}_{F}$ have dimension $4$ and $8$ over $F$, respectively.
Let $D$ be an alternative division ring. Observe that in general the set of polynomials with coefficients in $D$ is not uniquely determined. One may define $D[\mathtt{q}]$ to be the set of those polynomials having coefficients in $D$ on the left of the variable $\mathtt{q}$ (or on the right, respectively) but then the product which gives to $D[\mathtt{q}]$ a structure of ring is not the pointwise one, so the classical evaluation map $e_{a}:D[\mathtt{q}]\to D$ at $a\in D$ defined as $e_{a}\left(\sum_{h}^{m}a_{h}\mathtt{q}^{h}\right):=\sum_{h}^{m}a_{h}a^{h}$ is not a ring homomorphism if $a\in D\setminus R$. On the other hand, one may define $D[\mathtt{q}]$ as the set of $L$-terms with coefficients in $D$ in the variable $\mathtt{q}$. In this case the pointwise sum and product endow $D[\mathtt{q}]$ of a ring structure and those operations behave well with the evaluation map $e_{a}:D[\mathtt{q}]\to D$ at $a\in D$ which is actually a ring homomorphism but in that case polynomials are way much more complicated. However, since elements of $R$ actually commute with every element of $D$, such a choice is unique when we define the ring of polynomials $R[\mathtt{q}]$ in the variable $\mathtt{q}$ with coefficients in the center $R$ of $D$.
We say that $D$ is algebraically closed if every nonconstant polynomial $p(\mathtt{q})\in R[\mathtt{q}]$ has a root in $D$. For instance, both $\mathbb{H}$ and $\mathbb{O}$ are algebraically closed alternative division rings.
Remark 1.3.
Recall that $R:=\{a\in D\,|\,(ab)c=a(bc)=(ca)b,\,\forall b,c\in D\}$, so it is an $L$-definable set. This means that the algebraically closure condition for a an alternative division ring $D$ defined above can be expressed by a countable set of $L$-sentences. Indeed, next $L$-sentence means that each polynomial $p(\mathtt{q}):=\sum_{h=1}^{d}a_{h}\mathtt{q}^{h}\in R[\mathtt{q}]$ of fixed degree $d\geq 1$ has a root in $D$:
$\forall\mathtt{a}_{0}\dots\forall\mathtt{a}_{d}\left(\Big{(}\mathtt{a}_{0},\dots,\mathtt{a}_{d}\in R\wedge(a_{d}\neq 0)\Big{)}\rightarrow\Big{(}\exists\mathtt{b}\Big{(}\sum_{h=1}^{d}\mathtt{a}_{h}\mathtt{b}^{h}=0\Big{)}\Big{)}\right).$
Now we are in position to define the classes of $L$-structures containing $\mathbb{H}$ and $\mathbb{O}$, respectively, we are interested in.
Definition 1.4.
Let $H$ be an $L$-structure. Denote by $R:=\{q\in H\,|\,\forall p(p\cdot q=q\cdot p)\}$ the center of $H$. We say that $H$ is an algebraically closed quaternion algebra, ACQ for short, if the following axioms are satisfied:
  • (H1111)
    $H$ is an alternative division ring with unity such that $1\neq 0$.
  • (H2222)
    $H$ has dimension $4$ as an $R$-vector space, that is, there are $i,j,k\in H$ so that $\{1,i,j,k\}$ is a basis of $H$ as a vector space over its center $R$. As a first-order $L$-formula it is described as:
    $\quad\exists i\exists j\exists k\big{(}$$\neg(\exists\lambda_{i}\exists\lambda_{j}\exists\lambda_{k}((\lambda_{i},\lambda_{j},\lambda_{k}\in R)\wedge 1=\lambda_{i}\cdot i+\lambda_{j}\cdot j+\lambda_{k}\cdot k))\wedge$
    $\neg(\exists\lambda_{1}\exists\lambda_{j}\exists\lambda_{k}((\lambda_{1},\lambda_{j},\lambda_{k}\in R)\wedge i=\lambda_{1}+\lambda_{j}\cdot j+\lambda_{k}\cdot k))\wedge$
    $\neg(\exists\lambda_{1}\exists\lambda_{i}\exists\lambda_{k}((\lambda_{1},\lambda_{i},\lambda_{k}\in R)\wedge j=\lambda_{1}+\lambda_{i}\cdot i+\lambda_{k}\cdot k))\wedge$
    $\neg(\exists\lambda_{1}\exists\lambda_{i}\exists\lambda_{j}((\lambda_{1},\lambda_{i},\lambda_{j}\in R)\wedge k=\lambda_{1}+\lambda_{i}\cdot i+\lambda_{j}\cdot j))\wedge$
    $\forall q\exists\lambda_{1}\exists\lambda_{i}\exists\lambda_{j}\exists\lambda_{k}((\lambda_{1},\lambda_{i},\lambda_{j},\lambda_{k}\in R)\wedge(q=\lambda_{1}+\lambda_{i}\cdot i+\lambda_{j}\cdot j+\lambda_{k}\cdot k))\big{)}.$
  • (H3333)
    $H$ is algebraically closed.
Definition 1.5.
Let $O$ be an $L$-structure. Denote by $R:=\{o\in O\,|\,\forall a\forall b((oa)b=o(ab)=(bo)a)\}$ the center of $O$. We say that $O$ is an algebraically closed octonion algebra, ACO for short, if the following axioms are satisfied:
  • (O1111)
    $O$ is an alternative division ring with unity such that $1\neq 0$.
  • (O2222)
    $O$ has dimension $8$ as an $R$-vector space, that is, there are $e_{1},e_{2},e_{3},e_{4},e_{5},e_{6},$ $e_{7}\in O$ so that $\{1,e_{1},e_{2},e_{3},e_{4},e_{5},e_{6},e_{7}\}$ is a basis of $O$ as a vector space over its center $R$. As in the case of axiom $(\textnormal{H}2)$, previous property can be expressed as a first-order $L$-formula.
  • (O3333)
    $O$ is algebraically closed.
Remark 1.6.
Observe that, in general, to define a notion of vector space over a field a richer language is needed. Indeed, in general, one needs to define a functional symbol for each scalar multiplication with respect to an element of the ground field. However, in our case the language of rings $L:=\{+,-,\cdot,0,1\}$ is sufficient since the field $R$ such that $H$ and $O$ are $R$-vector spaces is a $L$-definable subset both in $H$ and $O$.
Denote by RCF the first-order theory of real closed fields. As a consequence of [Niv41, Theorem 1] and [Ghi12, Theorem 1.2] we have the following characterization of algebraically closed quaternion and octonion algebras:
Theorem 1.7.
Let $H$ and $O$ be alternative division rings and denote by $R$ their centers. Then:
  1. (i)
    $H\models\textnormal{ACQ}$ if and only if $R\models\textnormal{RCF}$ and $H$ is isomorphic to the quaternion algebra $\mathbb{H}_{R}$ over $R$.
  2. (ii)
    $O\models\textnormal{ACO}$ if and only if $R\models\textnormal{RCF}$ and $O$ is isomorphic to the octonion algebra $\mathbb{O}_{R}$ over $R$.
Let us specify previous theorem in terms of extension of $L$-substructures. Let $H_{1},H_{2}\models\textnormal{ACQ}$ and $O_{1},O_{2}\models\textnormal{ACO}$. We denote, with abuse of notation, $H_{1}$ (resp. $O_{1}$) is a $L$-substructure of $H_{2}$ (resp. $O_{2}$) as $H_{1}\subseteq H_{2}$ (resp. $O_{1}\subseteq O_{2}$). Similarly, we denote $H_{1}$ (resp. $O_{1}$) is an elementary L-substructure of $H_{2}$ (resp. $O_{2}$) as $H_{1}\preceq H_{2}$ (resp. $O_{1}\preceq O_{2}$).
Lemma 1.8.
Let $H\models\textnormal{ACQ}$ and $O\models\textnormal{ACO}$. Denote by $R$ the center of $H$ and $O$. Then, for every $R_{1}\models RCF$ such that $R_{1}\preceq R$, $\mathbb{H}_{R_{1}}$ can be embedded as a $L$-substructure of $H$ and $\mathbb{O}_{R_{1}}$ can be embedded as a $L$-substructure of $O$.
Proof.
Let $\mathbb{H}_{R_{1}}\subseteq\mathbb{H}_{R}$ and $\mathbb{O}_{R_{1}}\subseteq\mathbb{O}_{R}$. By Theorem 1.7, there are isomorphisms $\varphi_{H}:H\to\mathbb{H}_{R}$ and $\varphi_{O}:O\to\mathbb{O}_{R}$ so just consider the embeddings $\varphi_{H}^{-1}|_{\mathbb{H}_{R_{1}}}$ and $\varphi_{O}^{-1}|_{\mathbb{O}_{R_{1}}}$. ∎
Remark 1.9.
Denote by ¯$\overline{{\mathbb{Q}}}^{r}:=\overline{{\mathbb{Q}}}\cap{\mathbb{R}}$ the real closure of ${\mathbb{Q}}$. As a consequence of Lemma 1.8, we have that ¯$\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}$ can be embedded as a $L$-substructure of $H$ and $\mathbb{O}_{\overline{{\mathbb{Q}}}^{r}}$ can be embedded as a $L$-substructure of $O$, for every $H\models\textnormal{ACQ}$ and $O\models\textnormal{ACO}$.
Lemma 1.10.
Let $H_{1},H_{2}\models\textnormal{ACQ}$ and $O_{1},O_{2}\models\textnormal{ACO}$. Denote by $R_{1}$ the center of $H_{1}$ and $O_{1}$ and $R_{2}$ the center of $H_{2}$ and $O_{2}$. Then:
  1. (i)
    if $H_{1}\subseteq H_{2}$ and $\varphi_{1}:H_{1}\to\mathbb{H}_{R_{1}}$ is an isomorphism, then $\varphi_{1}$ extends uniquely to an isomorphism $\varphi_{2}:H_{2}\to\mathbb{H}_{R_{2}}$. In particular, $R_{1}\preceq R_{2}$.
  2. (ii)
    if $O_{1}\subseteq O_{2}$ and $\varphi_{1}:O_{1}\to\mathbb{O}_{R_{1}}$ is an isomorphism, then $\varphi_{1}$ extends uniquely to an isomorphism $\varphi_{2}:O_{2}\to\mathbb{O}_{R_{2}}$. In particular, $R_{1}\preceq R_{2}$.
Proof.
Let us first prove (i). Denote by $1_{H}$ the identity of $H_{1}$ and $H_{2}$ and by $i_{H}:=\varphi_{1}^{-1}(i),j_{H}:=\varphi_{1}^{-1}(j),k_{H}:=\varphi_{1}^{-1}(k)\in H_{1}\subset H_{2}$. Observe that $i_{H},j_{H},k_{H}\in H_{1}$ satisfy Definition 1.4$(\textnormal{H}2)$ for $H_{1}$. We will prove that $R_{1}\preceq R_{2}$ and $i_{H},j_{H},k_{H}\in H_{1}$ satisfy Definition 1.4$(\textnormal{H}2)$ for $H_{2}$ as well. Let $F:=R_{2}\cap H_{1}$. By definition of center of a ring, $F$ is an ordered subfield of $R_{1}$. Denote by $R:=\overline{F}^{r}$ the real closure of $F$, then $R\preceq R_{1}$. In addition, [BCR98, Theorem 1.2.2] ensures that $R\preceq R_{2}$ as well. By Lemma 1.8, there exists $H\subset H_{1}$ such that $H$ is isomorphic to $\mathbb{H}_{R}$, $i_{H},j_{H},k_{H}\in H$ and they satisfy Definition 1.4$(\textnormal{H}2)$ for $H$. Now, let us consider $H\subseteq H_{2}$. Let $i_{H}\in H$. Observe that, since $i_{H}\notin R$ but it is algebraic over $R$ and $R\preceq R_{2}$, we have that $i_{H}\notin R_{2}$. So $R[i_{H}]$ and $R_{1}[i_{H}]$ are algebraically closed fields of characteristic 00 such that $R[i_{H}]\preceq R_{1}[i_{H}]$ as $L$-structures in theory of ACF. In addition, $H=R[i_{H}]\oplus j_{H}R[i_{H}]$ is isomorphic to $\mathbb{H}_{R}$. Observe that $j_{H},k_{H}:=i_{H}j_{H}\notin R_{1}[i_{H}]$, otherwise $R_{1}[i_{H}]$ would not be commutative. Hence, as vector spaces over the real closed field $R_{2}$, we have $H_{2}=R[i_{H}]\oplus j_{H}R[i_{H}]$ as well. This proves that the map $\varphi_{2}:H_{2}\to\mathbb{H}_{R_{2}}$ defined by $R_{2}$-linear combinations of $\varphi_{2}(1_{H})=1$, $\varphi_{2}(i_{H})=i$, $\varphi_{2}(j_{H})=j$ and $\varphi_{2}(k_{H})=k$ is the unique isomorphism extending $\varphi_{1}$. The strategy to prove (ii) is the same as in (i). Denote by $1_{O}$ the identity of $O_{1}$ and $O_{2}$ and by $e_{\ell,O}:=\varphi_{1}^{-1}(e_{\ell})\in O_{1}\subset O_{2}$, for every $\ell=1,\dots,7$, where $\{1,e_{1},\dots,e_{7}\}$ denotes the canonical basis of $\mathbb{O}_{R_{1}}$. Define $F:=R_{2}\cap O_{1}$, it is a field so let $R:=\overline{F}^{r}$ be its real closure. Then, $R\preceq R_{2}$ as well. Observe that $\{e_{1,O},\dots,e_{7,O}\}\in O_{1}$ satisfy Definition 1.5$(\textnormal{O}2)$ for $O_{1}$. Let $O\subset O_{1}$ so that $O$ is isomorphic to $\mathbb{O}_{R}$, $e_{\ell,O}\in O$ for every $\ell=1,\dots,7$ and $\{e_{1,O},\dots,e_{7,O}\}$ satisfy Definition 1.5$(\textnormal{O}2)$ for $O$. Recall that $\mathbb{O}_{R}=\mathbb{H}_{R}\oplus e_{4}\mathbb{H}_{R}$, in which $i=e_{1,O}$, $j=e_{2,O}$ and $k=e_{3,O}$. So follow the same procedure of (i) getting a real algebra $H_{2}=R[e_{1,O}]\oplus e_{2,O}R[e_{1,O}]\subset O_{2}$ isomorphic to $\mathbb{H}_{R_{2}}$. Observe that $e_{4,O}\notin H_{2}$, otherwise $H_{2}$ would not be associative. Then, as a vector space over $R_{2}$, we have $O_{2}=H_{2}\oplus e_{4,H}H_{2}$. This proves that the map $\varphi_{2}:O_{2}\to\mathbb{O}_{R_{2}}$ defined by $R_{2}$-linear combinations of $\varphi_{2}(1_{O})=1$ and $\varphi_{2}(e_{\ell,O})=e_{\ell}$ for $\ell=1,\dots,7$ is the unique isomorphism extending $\varphi_{1}$. ∎
As a consequence of Theorem 1.7, up to isomorphism, we may suppose every model of ACQ is of the form $\mathbb{H}_{R}$ and every model of ACO is of the form $\mathbb{O}_{R}$, for some real closed field $R$. In addition, by Lemma 1.10 we may suppose for every $H_{1},H_{2}\models\textnormal{ACQ}$ and $O_{1},O_{2}\models\textnormal{ACO}$ such that $H_{1}\subseteq H_{2}$ and $O_{1}\subseteq O_{2}$ there exists $R_{1},R_{2}\models\textnormal{RCF}$ such that $H_{1}=\mathbb{H}_{R_{1}}\subseteq\mathbb{H}_{R_{2}}=H_{2}$ and $O_{1}=\mathbb{O}_{R_{1}}\subseteq\mathbb{O}_{R_{2}}=O_{2}$. So, from now on we will only deal with quaternion and octonion algebras over real closed fields as models of ACQ and ACO, respectively.
2. Model theoretical properties of ACQ &\&& ACO
This section is devoted to establish model theoretical properties of ACQ and ACO making use of (real) coordinates describing each model of these theories as a vector space over its center.
Consider $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$, for some $R\models\textnormal{RCF}$. Define the isomorphism of vector spaces $\varphi_{H}:\mathbb{H}_{R}\to R^{4}$ and $\varphi_{O};\mathbb{O}_{R}\to R^{8}$ sending each element of $H_{R}$ and $O_{R}$, respectively, to the string of its real coordinates. Namely: write uniquely each $q\in\mathbb{H}_{R}$ and $o\in\mathbb{O}_{R}$ as $q=x_{0}+ix_{1}+jx_{2}+kx_{3}$ and $o=x_{o}+\sum_{i=1}^{7}x_{i}e_{i}$, thus define $\varphi_{H}(q):=(x_{0},x_{1},x_{2},x_{3})$ and $\varphi_{O}(o):=(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})$.
Lemma 2.1.
Let $R\models\textnormal{RCF}$. The following hold:
  1. (i)
    Consider $\mathbb{H}_{R}$ and let $t(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ be a $L$-term. Write each variable in real coordinates as $\mathtt{q}_{\ell}:=\mathtt{x}_{\ell,0}+i\mathtt{x}_{\ell,1}+j\mathtt{x}_{\ell,2}+k\mathtt{x}_{\ell,3}$. Then there exists a $L$-boolean formula
    $\mathcal{B}(\mathtt{x}_{1,0},\mathtt{x}_{1,1},\mathtt{x}_{1,2},\mathtt{x}_{1,3},\dots,\mathtt{x}_{n,0},\mathtt{x}_{n,1},\mathtt{x}_{n,2},\mathtt{x}_{n,3})$
    such that for every $(q_{1},\dots,q_{n})\in\mathbb{H}_{R}^{n}$: $\mathbb{H}_{R}\models\big{(}t(q_{1},\dots,q_{n})=0\big{)}$ if and only if $R\models\mathcal{B}(x^{1}_{1},x^{2}_{1},x^{3}_{1},x^{4}_{1},\dots,x^{1}_{n},x^{2}_{n},x^{3}_{n},x^{4}_{n})$, with ${q}_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$, for every $\ell\in\{1,\dots,n\}$.
  2. (ii)
    Consider $\mathbb{O}_{R}$ and let $t(\mathtt{o}_{1},\dots,\mathtt{o}_{n})$ be a $L$-term. Write each variable in real coordinates as $\mathtt{o}_{\ell}:=\mathtt{x}_{\ell,0}+\sum_{h=1}^{7}\mathtt{x}_{\ell,h}e_{\ell}$. Then there exists a $L$-boolean formula
    $\mathcal{B}(\mathtt{x}_{1,0},\mathtt{x}_{1,1},\dots,\mathtt{x}_{1,7},\dots,\mathtt{x}_{n,0},\mathtt{x}_{n,1},\dots,\mathtt{x}_{n,7})$
    such that for every $(o_{1},\dots,o_{n})\in\mathbb{O}_{R}^{n}$: $\mathbb{O}_{R}\models\big{(}t(q_{1},\dots,q_{n})=0\big{)}$ if and only if $R\models\mathcal{B}(x_{1,0},x_{1,1},\dots,x_{1,7},\dots,x_{n,0},x_{n,1},\dots,x_{n,7})$ with ${o}_{\ell}:=x_{\ell,0}+\sum_{h=1}^{7}x_{\ell,h}e_{\ell}$, for every $\ell\in\{1,\dots,n\}$.
Proof.
We only prove (i) since the same strategy works also for (ii). Observe that, since $\mathbb{H}_{R}\models\textnormal{ACQ}$ is not commutative and $L$ is the language of rings, $L$-terms depending on $n$-variables just consist of polynomials with integer coefficients in the non-commutating variables $\mathtt{q}_{1},\dots,\mathtt{q}_{n}$. Namely, each polynomial consists of a finite sum of monomials of the form $\prod_{h=1}^{m}\mathtt{q}_{i_{h}}^{s_{h}}$, with $i_{h}\in\{1,\dots,n\}$ and $i_{h}\neq i_{h+1}$, for every $h\in\{1,\dots,m-1\}$. Let $\{1,i,j,k\}$ be the standard basis of $\mathbb{H}_{R}$ as a vector space over $R$. Substitute each $\mathtt{q}_{\ell}$ with $\mathtt{x}_{\ell,0}+i\mathtt{x}_{\ell,1}+j\mathtt{x}_{\ell,2}+k\mathtt{x}_{\ell,3}$. Recall that previous scripture is unique when we only admit real values for variables $\mathtt{x}_{\ell,0},\mathtt{x}_{\ell,1},\mathtt{x}_{\ell,2}$ and $\mathtt{x}_{\ell,3}$. Thus, unfold the computation isolating the elements $i,j,k$ assuming that variables $\{\mathtt{x}_{\ell,0},\mathtt{x}_{\ell,1},\mathtt{x}_{\ell,2},\mathtt{x}_{\ell,3}\}_{\ell=1}^{n}$ commute each other and with elements $i,j,k\in\mathbb{H}_{R}$. Thus, there are unique polynomials $p_{1},p_{2},p_{3},p_{4}\in{\mathbb{Z}}[\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3}]$ such that $t(q_{1},\dots,q_{n})=0$ if and only if $p_{1}(x_{1,0},\dots,x_{n,3})+ip_{2}(x_{1,0},\dots,x_{n,3})+jp_{3}(x_{1,0},\dots,x_{n,3})+kp_{4}(x_{1,0},\dots,x_{n,3})=0$, with $q_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$ for unique $x_{\ell,0},x_{\ell,1},x_{\ell,2},x_{\ell,3}\in R$, for every $\ell\in\{1,\dots,n\}$. Then, since ${1,i,j,k}$ is a basis of $\mathbb{H}_{R}$ as a vector space over $R$, it follows that $\mathbb{H}_{R}\models t(q_{1},\dots,q_{n})=0$ if and only if $R\models\bigwedge_{s=1}^{4}(p_{s}(x{1,0},\dots,x_{n,3})=0)$. Hence, just take
$\mathcal{B}(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3}):=\bigwedge_{s=1}^{4}\big{(}p_{s}(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3})=0\big{)}.$
As a corollary we obtain a way of tranlating first-order formulas $\phi$ for quaternion and octonion algebras over real closed fields to formulas in real coordinates.
Corollary 2.2.
Let $R\models\textnormal{RCF}$. The following hold:
  1. (i)
    Let $\phi(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ be a first-order $L$-formula. Write each variable in real coordinates as $\mathtt{q}_{\ell}:=\mathtt{x}_{\ell,0}+i\mathtt{x}_{\ell,1}+j\mathtt{x}_{\ell,2}+k\mathtt{x}_{\ell,3}$. Then, there exists a first-order $L$-formula
    $\psi(\mathtt{x}_{1,0},\mathtt{x}_{1,1},\mathtt{x}_{1,2},\mathtt{x}_{1,3},\dots,\mathtt{x}_{n,0},\mathtt{x}_{n,1},\mathtt{x}_{n,2},\mathtt{x}_{n,3})$
    such that for every $(q_{1},\dots,q_{n})\in\mathbb{H}_{R}^{n}$: $\mathbb{H}_{R}\models\phi(q_{1},\dots,q_{n})$ if and only if $R\models\psi(x_{1,0},x_{1,1},x_{1,2},x_{1,3},\dots,x_{n,0},x_{n,1},x_{n,2},x_{n,3})$, with ${q}_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$, for every $\ell\in\{1,\dots,n\}$.
  2. (ii)
    Let $\phi(\mathtt{o}_{1},\dots,\mathtt{o}_{n})$ be a first-order $L$-formula. Write each variable in real coordinates as $\mathtt{o}_{\ell}:=\mathtt{x}_{\ell,0}+\sum_{h=1}^{7}\mathtt{x}_{\ell,h}e_{\ell}$. Then, there exists a first-order $L$-formula
    $\psi(\mathtt{x}_{1,0},\mathtt{x}_{1,1},\dots,\mathtt{x}_{1,7},\dots,\mathtt{x}_{n,0},\mathtt{x}_{n,1},\dots,\mathtt{x}_{n,7})$
    such that for every $(o_{1},\dots,o_{n})\in\mathbb{O}_{R}^{n}$: $\mathbb{O}_{R}\models\phi(o_{1},\dots,o_{n})$ if and only if $R\models\psi(x_{1,0},x_{1,1},\dots,x_{1,7},\dots,x_{n,0},x_{n,1},\dots,x_{n,7})$, with ${o}_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$, for every $\ell\in\{1,\dots,n\}$.
Proof.
By Lemma 2.1 and induction on $L$-formulas. ∎
Remark 2.3.
Observe that if $\varphi$ of Corollary 2.2 is boolean, then the resulting $\psi$ is boolean as well. In addition, Corollary 2.2 holds true if we consider the languages $L_{R}:=L\cup\{c_{i}\}_{i\in R}$, $L_{\mathbb{H}_{R}}:=L\cup\{q_{i}\}_{i\in\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}:=L\cup\{o_{i}\}_{i\in\mathbb{O}_{R}}$ be the extensions of $L$ with constant symbols from $R$, $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$, respectively.
2.1. Model Completeness &\&& Completeness
Here we are in position to prove the main properties of ACQ and ACO contained in this paper.
Theorem 2.4.
The $L$-theories of ACQ and ACO are model complete.
Proof.
We only prove the case of ACQ, the proof is exactly the same for ACO. Let $R_{1},R_{2}\models\textnormal{RCF}$ such that $R_{1}\preceq R_{2}$ and consider $\mathbb{H}_{R_{1}}\subseteq\mathbb{H}_{R_{2}}$. We prove $\mathbb{H}_{R_{1}}\preceq\mathbb{H}_{R_{2}}$. Let $\phi(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ and $q_{1},\dots,q_{n}\in\mathbb{H}_{R_{1}}$. By Corollary 2.2 there exists a formula $\psi(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3})$ such that: $\mathbb{H}_{R_{1}}\models\phi(q_{1},\dots,q_{n})$ if and only if $R_{1}\models\psi(x_{1,0},\dots,x_{n,3})$, with $x_{\ell,0},x_{\ell,1},x_{\ell,2},x_{\ell,3}\in R_{1}$ such that $q_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$, for every $\ell\in\{1,\dots,n\}$. By model completeness of the $L$-theory of real closed fields we have that $R_{1}\models\psi(x_{1,0},\dots,x_{n,3})$ if and only if $R_{2}\models\psi(x_{1,0},\dots,x_{n,3})$. Since $\mathbb{H}_{R_{1}}\subseteq\mathbb{H}_{R_{2}}$ and elements $x_{\ell,0},x_{\ell,1},x_{\ell,2},x_{\ell,3}\in R_{1}$ such that $q_{\ell}=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$ are unique in $R_{2}$ too for every $\ell\in\{1,\dots,n\}$, since $\{1,i,j,k\}$ is a basis of $\mathbb{H}_{R_{2}}$ as a vector space over $R_{2}$, we conclude that: $R_{2}\models\psi(x_{1,0},\dots,x_{n,3})$ if and only if $\mathbb{H}_{R_{2}}\models\phi(q_{1},\dots,q_{n})$, as desired. ∎
Theorem 2.5.
The $L$-theories of ACQ and ACO are complete. In particular, they coincide with $\textnormal{Th}_{L}(\mathbb{H})$ and $\textnormal{Th}_{L}(\mathbb{O})$, respectively.
Proof.
Let $R\models\textnormal{RCF}$. Recall that ¯$\overline{{\mathbb{Q}}}^{r}$, namely the real closure of ${\mathbb{Q}}$, is the smallest real closed field, so ¯$\overline{{\mathbb{Q}}}^{r}\preceq R$, ¯$\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}\preceq\mathbb{H}_{R}$ and $\mathbb{O}_{\overline{{\mathbb{Q}}}^{r}}\preceq\mathbb{O}_{R}$. By model completeness of ACQ and ACO, namely Theorem 2.4, we deduce that both the $L$-theories ACQ and ACO are complete. ∎
2.2. Failure of Quantifier Elimination
Here we continue the comparison between $L$-formulas in $n$ quaternion or octonion variables with $L$-formulas in $2n$ or $4n$ real variables, respectively. Here we introduce a well-known result in semialgebraic geometry, that is: Every semialgebriac set $S\subset R^{n}$ is the image under projection of an algebraic set $V\subset R^{m}:=R^{n}\times R^{m-n}$. Fore sake of completeness we restate the result and we give a proof.
Lemma 2.6.
Let $R\models\textnormal{RCF}$ and let $L_{<}:=L\cup\{<\}$ be the language of ordered fields. Then, for every $L_{<}$-formula $\phi(\mathtt{x}_{1},\dots,\mathtt{x}_{n})$ there exists a $L$-formula $\psi(\mathtt{x}_{1},\dots,\mathtt{x}_{n})$ of the form
$\psi(\mathtt{x}_{1},\dots,\mathtt{x}_{n}):=\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{m}(p(\mathtt{x}_{1},\dots,\mathtt{x}_{n},\mathtt{y}_{1},\dots,\mathtt{x}_{m})),$
for some $m\in{\mathbb{N}}$ such that $R\models\phi(x_{1},\dots,x_{n})$ if and only if $R\models\psi(x_{1},\dots,x_{n})$ for every $x_{1},\dots,x_{n}\in R$.
Proof.
Denote by $\mathtt{x}:=(\mathtt{x}_{1},\dots,\mathtt{x}_{n})$ and by $x:=(x_{1},\dots,x_{n})\in R^{n}$. By quantifier elimination of the $L_{<}$-theory of real closed fields we may suppose $\phi(\mathtt{x})$ is a boolean formula, thus write $\phi(\mathtt{x})$ in disjunctive normal form, that is:
$\phi(\mathtt{x})\equiv\bigvee_{s=1}^{a}\big{(}(p_{s0}(\mathtt{x})=0)\wedge\bigwedge_{t=1}^{b_{s}}(0<p_{st}(\mathtt{x}))\big{)}.$
CLAIM: Each $L_{<}$-formula $\phi_{s}(\mathtt{x}):=(p_{s0}(\mathtt{x})=0)\wedge\bigwedge_{t=1}^{b_{s}}(0<p_{st}(\mathtt{x}))$ is equivalent to a $L$-formula of the form $\psi_{s}(\mathtt{x}):=\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s+1}}(p(\mathtt{x},\mathtt{y}_{1},\dots,\mathtt{y}_{m_{s}})=0)$. Consider the $L_{<}$-formula $\phi^{\prime}_{s}(\mathtt{x}):=\bigwedge_{t=1}^{b_{s}}(0<p_{st}(\mathtt{x})$. Observe that $(0<p_{st}(\mathtt{x}))\equiv\exists\mathtt{y}_{t}((\mathtt{y}_{t}^{2}=p_{st}(\mathtt{x}))\wedge(p_{st}(\mathtt{x})\neq 0)\big{)}$ for every $t=1,\dots,b_{s}$, thus also
$\phi^{\prime}_{s}(\mathtt{x})\equiv\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s}}\Big{(}\bigwedge_{t=1}^{b_{s}}\big{(}(\mathtt{y}_{t}^{2}=p_{st}(\mathtt{x}))\wedge(p_{st}(\mathtt{x})\neq 0)\big{)}\Big{)}.$
Let $r_{s}(\mathtt{x},\mathtt{y}):=\sum_{t=1}^{b_{s}}(\mathtt{y}_{t}^{2}-p_{st}(\mathtt{x}))^{2}$ and $q_{s}(\mathtt{x},\mathtt{y}):=r_{s}(\mathtt{x},\mathtt{y})^{2}+\prod_{t=1}^{b_{s}}p_{st}(\mathtt{x})^{2}$. Then:
$\phi^{\prime}_{s}(\mathtt{x})$ $\equiv\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s}}\exists\mathtt{z}(r_{s}(\mathtt{x})=0)\wedge(q_{s}(\mathtt{x})\mathtt{z}=1)$
$\equiv\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s}}\exists\mathtt{y}_{b_{s+1}}\big{(}r^{2}_{s}(\mathtt{x})+(q_{s}(\mathtt{x})\mathtt{y}_{b_{s+1}}-1)^{2}=0\big{)}.$
Let $p^{\prime}_{s0}(\mathtt{x},\mathtt{y}_{1},\dots,\mathtt{y}_{b_{s+1}}):=r^{2}_{s0}(\mathtt{x})+\sum_{t=1}^{b_{s}+1}\mathtt{y}_{t}^{2}$. Then:
$\phi_{s}(\mathtt{x})$ $\equiv(r^{2}_{s0}(\mathtt{x})=0)\wedge\phi^{\prime}_{s}(\mathtt{x})$
$\equiv(r^{2}_{s0}(\mathtt{x})=0)\wedge\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s}}\exists\mathtt{y}_{b_{s+1}}\big{(}r^{2}_{s}(\mathtt{x})+(q_{s}(\mathtt{x})\mathtt{y}_{b_{s+1}}-1)^{2}=0\big{)}$
$\equiv\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{b_{s}}\exists\mathtt{y}_{b_{s+1}}\big{(}r^{2}_{s}(\mathtt{x})+(q_{s}(\mathtt{x})\mathtt{y}_{b_{s+1}}-1)^{2}+p^{\prime}_{s0}(\mathtt{x},\mathtt{y}_{1},\dots,\mathtt{y}_{b_{s+1}})=0\big{)},$
so the CLAIM is proved. Let $p_{s}(\mathtt{x},\mathtt{y}_{s,1},\dots,\mathtt{y}_{s,b_{s}+1}):=r^{2}_{s}(\mathtt{x})+(q_{s}(\mathtt{x})\mathtt{y}_{s,b_{s+1}}-1)^{2}+p^{\prime}_{s0}(\mathtt{x},\mathtt{y}_{s,1},\dots,\mathtt{y}_{s,b_{s+1}})$ for every $s=1,\dots,a$. Let $b:=\max\{b_{s}\}_{s=1}^{a}$ and $\mathtt{y}:=(\mathtt{y}_{1},\dots,\mathtt{y}_{b+1})$. Define the polynomial $p^{\prime}_{s}(\mathtt{x},\mathtt{y})$ for every $s=1,\dots,a$ as follows:
$p^{\prime}_{s}(\mathtt{x},\mathtt{y}):=p_{s}(\mathtt{x},\mathtt{y}_{1},\dots,\mathtt{y}_{b_{s}+1})+\sum_{t=s+2}^{b+1}\mathtt{y}^{2}_{t}.$
Hence, we obtain that
$\phi(\mathtt{x})\equiv\exists\mathtt{x}\exists\mathtt{y}\big{(}\prod_{s=1}^{a}p^{\prime}_{s}(\mathtt{x},\mathtt{y})=0\big{)},$
as desired. ∎
Next result constitutes the reverse counterpart of Lemma 2.1
Lemma 2.7.
Let $R\models\textnormal{RCF}$ and denote by $L_{<}:=L\cup\{<\}$ the language of ordered fields.
  1. (i)
    Let $L^{\prime}:=L\cup\{i,j,k\}$ be the extension of $L$ with constant symbols $\{i,j,k\}$ whose interpretations are the elements $i,j,k\in\mathbb{H}_{R}$. Then, for every first-order $L_{<}$-formula $\phi(\mathtt{x}_{1,0},\mathtt{x}_{1,1},\mathtt{x}_{1,2},\mathtt{x}_{1,3},\dots,\mathtt{x}_{n,0},\mathtt{x}_{n,1},\mathtt{x}_{n,2},\mathtt{x}_{n,3})$ there exists a $L^{\prime}$-formula $\psi(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ such that, for every $(x_{1,0},\dots,x_{n,3})\in R^{4n}$, the following holds:
    $R\models\phi(x_{1,0},\dots,x_{n,3})\quad\text{if and only if}\quad\mathbb{H}_{R}\models\psi(q_{1},\dots,q_{n}).$ (1)
    with $q_{\ell}:=x_{\ell,0}+ix_{\ell,1}+jx_{\ell,2}+kx_{\ell,3}$ for every $\ell\in\{1,\dots,n\}$.
  2. (ii)
    Let $L^{\prime}:=L\cup\{e_{h}\}_{h=1}^{7}$ be the extension of $L$ with constant symbols $\{e_{h}\}_{h=1}^{7}$ whose interpretations are the elements $\{e_{h}\}_{h=1}^{7}\in\mathbb{O}_{R}$. Then, for every first-order $L_{<}$-formula $\phi(\mathtt{x}_{1,0},\dots,\mathtt{x}_{1,7},\dots,\mathtt{x}_{n,0},\dots,\mathtt{x}_{n,7})$ there exists a $L^{\prime}$-formula $\psi(\mathtt{o}_{1},\dots,\mathtt{o}_{n})$ such that, for every $(x_{1,0},\dots,x_{n,7})\in R^{8n}$, the following holds:
    $R\models\phi(x_{1,0},\dots,x_{n,7})\quad\text{if and only if}\quad\mathbb{O}_{R}\models\psi(o_{1},\dots,o_{n}).$ (2)
    with $o_{\ell}:=x_{\ell,0}+\sum_{h=1}^{7}e_{h}x_{\ell,h}$ for every $\ell\in\{1,\dots,n\}$.
Proof.
Let us prove (i). Let $\phi(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3})$ be a $L_{<}$-formula. By quantifier elimination of the $L_{<}$-theory of RCF and Lemma 2.6, we may suppose that $\phi$ is expressed in the language $L$ and is of the following form:
$\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{m}\Big{(}p(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3},\mathtt{y}_{1},\dots,\mathtt{y}_{m})=0\Big{)},$
for some $m\in{\mathbb{N}}$. Observe that we may suppose $4|m$, indeed if it was not the case we can consider $m^{\prime}$ to be the minimum multiple of $4$ greater than $m$ and then consider the $L$-formula:
$\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{m^{\prime}}\Big{(}p(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,3},\mathtt{y}_{1},\dots,\mathtt{y}_{m})^{2}+\sum_{s=m+1}^{m^{\prime}}\mathtt{y}_{s}^{2}=0\Big{)},$
So let $m=4n^{\prime}$, for some $n^{\prime}\in{\mathbb{N}}$. Define the polynomial $p^{\prime}(\mathtt{q_{1}},\dots,\mathtt{q}_{n},\mathtt{r}_{1},\dots,\mathtt{r}_{n^{\prime}})$ with coefficients in $\mathbb{H}_{R}$ by substituting each real variable of $p(\mathtt{x},\mathtt{y})$ with the following expressions:
$\mathtt{x}_{s,0}$ $=\frac{1}{4}(\mathtt{q}_{s}-i\mathtt{q}_{s}i-j\mathtt{q}_{s}j-k\mathtt{q}_{s}k),\quad\,\,\,\mathtt{x}_{s,1}=\frac{1}{4i}(\mathtt{q}_{s}-i\mathtt{q}_{s}i+j\mathtt{q}_{s}j+k\mathtt{q}_{s}k),$ (3)
$\mathtt{x}_{s,2}$ $=\frac{1}{4j}(\mathtt{q}_{s}+i\mathtt{q}_{s}i-j\mathtt{q}_{s}j+k\mathtt{q}_{s}k),\quad\mathtt{x}_{s,3}=\frac{1}{4k}(\mathtt{q}_{s}+i\mathtt{q}_{s}i+j\mathtt{q}_{s}j-k\mathtt{q}_{s}k),$ (4)
$\mathtt{y}_{t,0}$ $=\frac{1}{4}(\mathtt{r}_{t}-i\mathtt{r}_{t}i-j\mathtt{r}_{t}j-k\mathtt{r}_{t}k),\quad\,\,\,\,\,\mathtt{y}_{t,1}=\frac{1}{4i}(\mathtt{r}_{t}-i\mathtt{r}_{t}{i}+{j}\mathtt{r}_{t}{j}+{k}\mathtt{r}_{t}{k}),$ (5)
$\mathtt{y}_{t,2}$ $=\frac{1}{4{j}}(\mathtt{r}_{t}+{i}\mathtt{r}_{t}{i}-{j}\mathtt{r}_{t}{j}+{k}\mathtt{r}_{t}{k}),\quad\,\,\mathtt{y}_{t,3}=\frac{1}{4{k}}(\mathtt{r}_{t}+{i}\mathtt{r}_{t}{i}+{j}\mathtt{r}_{t}{j}-{k}\mathtt{r}_{t}{k}),$ (6)
for every $s\in\{1,\dots,n\}$ and $t\in\{1,\dots,n^{\prime}\}$. For every $(x,y):=(x_{1,0},\dots,x_{n,3},y_{1,0},\dots,y_{n^{\prime},3})\in R^{4n}\times R^{4n^{\prime}}$ define $(q,r):=(q_{1},\dots,q_{m},r_{1},\dots,r_{s})\in\mathbb{H}_{R}^{n}\times\mathbb{H}_{R}^{n^{\prime}}$ as $q_{s}:=x_{s,0}+ix_{s,1}+jx_{s,2}+kx_{s,3}$, for every $s\in\{1,\dots,n\}$, and $r_{t}:=y_{t,0}+iy_{t,1}+jy_{t,2}+ky_{t,3}$, for every $t\in\{1,\dots,n^{\prime}\}$. Then, we obtain that
$R\models\phi(x)\quad\text{if and only if}\quad\mathbb{H}_{R}\models\exists\mathtt{r}_{1}\dots\exists\mathtt{r}_{n^{\prime}}\big{(}p^{\prime}(q_{1},\dots,q_{n},\mathtt{r}_{1},\dots,\mathtt{r}_{n^{\prime}})=0\big{)},$
thus the thesis follows. To prove (ii) the strategy is similar. Let $\phi$ be of the form:
$\exists\mathtt{y}_{1}\dots\exists\mathtt{y}_{m}\Big{(}p(\mathtt{x}_{1,0},\dots,\mathtt{x}_{n,7},\mathtt{y}_{1},\dots,\mathtt{y}_{m})=0\Big{)},$
with $m=8n^{\prime}$ for some $n^{\prime}\in{\mathbb{N}}$. Define the polynomial $p^{\prime}(\mathtt{o_{1}},\dots,\mathtt{o}_{n},\mathtt{r}_{1},\dots,\mathtt{r}_{n^{\prime}})$ with coefficients in $\mathbb{O}_{R}$ by substituting each real variable of $p(\mathtt{x},\mathtt{y})$ with the following expressions:
$\mathtt{x}_{s,0}$ $=\frac{1}{8}\Big{(}\mathtt{o}_{s}-\sum_{h=1}^{7}e_{h}\mathtt{o}_{s}e_{h}\Big{)},\quad\,\,\,\mathtt{x}_{s,\ell}=\frac{1}{8e_{\ell}}\Big{(}\mathtt{o}_{s}-e_{\ell}\mathtt{o}_{s}e_{\ell}+\sum_{h\neq\ell}e_{h}\mathtt{o}_{s}e_{h}\Big{)},$ (7)
$\mathtt{y}_{t,0}$ $=\frac{1}{8}\Big{(}\mathtt{r}_{t}-\sum_{h=1}^{7}e_{h}\mathtt{q}_{s}e_{h}\Big{)},\quad\,\,\,\mathtt{y}_{t,\ell}=\frac{1}{8e_{\ell}}\Big{(}\mathtt{r}_{t}-e_{\ell}\mathtt{r}_{t}e_{\ell}+\sum_{h\neq\ell}e_{h}\mathtt{r}_{t}e_{h}\Big{)},$ (8)
for every $s\in\{1,\dots,n\}$, $t\in\{1,\dots,n^{\prime}\}$ and $\ell\in\{1,\dots,7\}$. For every $(x,y):=(x_{1,0},\dots,x_{n,7},y_{1,0},\dots,y_{n^{\prime},7})\in R^{8n}\times R^{8n^{\prime}}$ define $(o,r):=(o_{1},\dots,o_{m},r_{1},\dots,r_{s})\in\mathbb{O}_{R}^{n}\times\mathbb{O}_{R}^{n^{\prime}}$ as $o_{s}:=x_{s,0}+\sum_{h=1}^{7}e_{h}x_{s,h}$, for every $s\in\{1,\dots,n\}$, and $r_{t}:=y_{t,0}+\sum_{h=1}^{7}e_{h}y_{s,h}$, for every $t\in\{1,\dots,n^{\prime}\}$. Then, we obtain that
$R\models\phi(x)\quad\text{if and only if}\quad\mathbb{O}_{R}\models\exists\mathtt{r}_{1}\dots\exists\mathtt{r}_{n^{\prime}}\big{(}p^{\prime}(o_{1},\dots,o_{n},\mathtt{r}_{1},\dots,\mathtt{r}_{n^{\prime}})=0\big{)},$
thus the thesis follows. ∎
Remark 2.8.
Observe that the elements $i,j,k\in\mathbb{H}_{R}$ and $\{e_{h}\}_{h=1,\dots,7}\in\mathbb{O}_{R}$ are not $L$-definable. Let us focus on the quaternionic case, a similar argument works for octonions as well. Suppose $i$ is a $L$-definable element, then restricting the $L$-formula defining $i$ to the complex plane $C=R\oplus Ri$ gives, as a consequence, that $i$ is $L$-definable in $C$ as well but this is impossible by unique factorization of polynomials over ${\mathbb{Q}}$. This proves that it was necessary to add at least constants $i,j,k\in\mathbb{H}_{R}$ and $\{e_{h}\}_{h=1,\dots,7}\in\mathbb{O}_{R}$ to the language $L$ to obtain equations (4),(5)&(8). We point out that formulas (4) and (5) where already known in literature, see [GS12] as a reference, whereas formulas (8) seem to appear for the first time even though they derive from a simple computation similar to the quaternionionic case.
Remark 2.9.
Fix $R\models\textnormal{RCF}$. Observe that Lemmas 2.6&2.7 hold true as well considering the languages $L_{<,R}:=L_{<}\cup\{c_{i}\}_{i\in R}$, that is, the extension of $L_{<}$ with constant symbols from $R$, and the languages $L_{\mathbb{H}_{R}}:=L\cup\{q_{i}\}_{i\in\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}:=L\cup\{o_{i}\}_{i\in\mathbb{O}_{R}}$, that are, the extensions of $L$ with constant symbols from $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$, respectively. We observe that considering latter languages Lemmas 2.6&2.7 provide a proof of model completeness for the theories of quaternions and octonions with respect to mentioned languages, see [Hod93, Theorem 8.3.1.].
Remark 2.10.
Consider the languages $L_{\mathbb{H}_{R}}:=L\cup\{q_{i}\}_{i\in\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}:=L\cup\{o_{i}\}_{i\in\mathbb{O}_{R}}$ as above. Consider the conjugation function :${}^{c}_{\mathbb{H}}:\mathbb{H}\to\mathbb{H}$ and :${}^{c}_{\mathbb{O}}:\mathbb{O}\to\mathbb{O}$ defined as $q^{c}_{\mathbb{H}}:=a_{0}-a_{1}i-a_{2}j-a_{3}k$ and $o^{c}_{\mathbb{H}}:=a_{0}-\sum_{h=1}^{7}a_{h}e_{h}$. Then, by Remark 2.9, the functions ${}^{c}_{\mathbb{H}}$ and ${}^{c}_{\mathbb{O}}$ are $L_{\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}$-definable, respectively.
Here we prove that both ACQ and ACO do not admit quantifier elimination.
Denote by $L_{R}:=L\cup\{c_{i}\}_{i\in R}$, $L_{\mathbb{H}_{R}}:=L\cup\{q_{i}\}_{i\in\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}:=L\cup\{o_{i}\}_{i\in\mathbb{O}_{R}}$ be the extensions of $L$ with coefficients in $R$, $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$, respectively.
Theorem 2.11.
Let $R\models\textnormal{RCF}$. Denote by $\textnormal{Th}_{L_{\mathbb{H}_{R}}}(\mathbb{H}_{R})$ the $L_{\mathbb{H}_{R}}$-theory of $\mathbb{H}_{R}$ and by $\textnormal{Th}_{L_{\mathbb{O}_{R}}}(\mathbb{O}_{R})$ the $L_{\mathbb{O}_{R}}$-theory of $\mathbb{O}_{R}$. Then, $\textnormal{Th}_{L_{\mathbb{H}_{R}}}(\mathbb{H}_{R})$ and $\textnormal{Th}_{L_{\mathbb{O}_{R}}}(\mathbb{O}_{R})$ do not have quantifier elimination.
Proof.
We just prove the statement for $\textnormal{Th}_{L_{\mathbb{H}_{R}}}(\mathbb{H}_{R})$ since the strategy is exactly the same for $\textnormal{Th}_{L_{\mathbb{O}_{R}}}(\mathbb{O}_{R})$. Consider the formula defining the unitary ball in $\mathbb{H}_{R}$, that is consider the $L_{<}$-formula $\phi(\mathtt{x_{0}},\dots,\mathtt{x}_{3}):=\big{(}\sum_{h=0}^{3}x_{h}^{2}\leq 1\big{)}$. Denote by $B(0,1):=\{(x_{0},\dots,x_{3})\in R^{4}\,|\,\phi(x_{0},\dots,x_{3})\}$ the unitary ball in $R^{4}$. Since boolean $L_{R}$-formulas define exactly subsets of $R^{4}$ which are Zariski open in their Zariski closure and every polynomial $p(\mathtt{x}_{0},\dots,\mathtt{x}_{3})\in R[\mathtt{x}_{0},\dots,\mathtt{x}_{3}]$ vanishing either over $B(0,1)$ or over $R^{4}\setminus B(0,1)$ is forced to vanish over the whole $R^{4}$, we deduce that $\phi(\mathtt{x_{0}},\dots,\mathtt{x}_{3})$ is not equivalent to any boolean $L_{R}$-formula. Hence, by Lemma 2.7, there exists a $L_{\mathbb{H}_{R}}$-formula $\psi(\mathtt{q})$ such that: $R\models\phi(x_{0},\dots,x_{3})$ if and only if $\mathbb{H}_{R}\models\psi(q)$, for every $x_{0},\dots,x_{3}\in R$ with $q:=x_{0}+ix_{1}+jx_{2}+kx_{3}\in\mathbb{H}_{R}$. However, by Remark 2.3, we have that $\psi(\mathtt{q})$ is equivalent to some boolean $L_{\mathbb{H}_{R}}$-formula if and only if $\phi(\mathtt{x_{0}},\dots,\mathtt{x_{3}})$ is equivalent to some boolean $L_{R}$-formula. Since $\phi(\mathtt{x_{0}},\dots,\mathtt{x_{3}})$ is not, we conclude that $\psi(\mathtt{q})$ is not equivalent to any boolean $L_{\mathbb{H}_{R}}$-formula, as desired. ∎
Hence, as a consequence we get the following general result.
Corollary 2.12.
ACQ and ACO do not have quantifier elimination.
Proof.
Suppose ACQ and ACO do have quantifier elimination. Let $R\models\textnormal{RCF}$. Then, a fortiori, $\textnormal{Th}_{L_{\mathbb{H}_{R}}}(\mathbb{H}_{R})$ and $\textnormal{Th}_{L_{\mathbb{O}_{R}}}(\mathbb{O}_{R})$ would have quantifier elimination, but this contradicts Theorem 2.11. ∎
3. Ordered polynomials: the fragments of ordered formulas
3.1. Algebraic sets & real dimension
In this section we will consider a fragment of the theories ACQ and ACO. Let us introduce the set of $L$-terms we are admitting in this fragment.
Definition 3.1.
Let $P$ be a $L$-polynomial in the variables $\mathtt{q}_{1},\dots,\mathtt{q}_{n}$. We say that $P$ is a ordered $L$-polynomial if it is the finite sum of monomials of the form
$\mathtt{q}_{1}^{\alpha_{1}}\cdot\mathtt{q}_{2}^{\alpha_{2}}\cdot\,\dots\,\cdot\mathtt{q}_{n}^{\alpha_{n}},$
for some $\alpha:=(\alpha_{1},\dots,\alpha_{n})\in{\mathbb{N}}^{n}$. We refer to the fragment of ordered $L$-formulas as the fragment of ACQ or ACO, respectively, in which the admitted $L$-terms are exactly ordered $L$-polynomials.
Observe that Definition 3.1 extends naturally when we admit coefficients in some fixed structure of ACQ or ACO.
Definition 3.2.
Let $R\models\textnormal{RCF}$. Let $P$ be a $L_{\mathbb{H}_{R}}$-polynomial or a $L_{\mathbb{O}_{R}}$-polynomial in the variables $\mathtt{q}_{1},\dots,\mathtt{q}_{n}$. We say that $P$ is a ordered polynomial with coefficients in $\mathbb{H}_{R}$ or in $\mathbb{O}_{R}$, respectively, if it is the finite sum of monomials of the form
$\mathtt{q}_{1}^{\alpha_{1}}\cdot\mathtt{q}_{2}^{\alpha_{2}}\cdot\,\dots\,\cdot\mathtt{q}_{n}^{\alpha_{n}}a_{\alpha},$
for some $\alpha:=(\alpha_{1},\dots,\alpha_{n})\in{\mathbb{N}}^{n}$ and $a_{\alpha}\in\mathbb{H}_{R}$ or $a_{\alpha}\in\mathbb{O}_{R}$, respectively. We refer to the fragment of ordered formulas with coefficients in $\mathbb{H}_{R}$ or in $\mathbb{O}_{R}$, respectively, as the fragments of $L_{\mathbb{H}_{R}}$-formulas or $L_{\mathbb{O}_{R}}$-formulas in which the admitted $L_{\mathbb{H}_{R}}$-terms or the admitted $L_{\mathbb{O}_{R}}$-terms are exactly ordered polynomials with coefficients in $\mathbb{H}_{R}$ or $\mathbb{O}_{R}$, respectively.
There are at least two motivations to be interested in studying the fragment of ordered $L$-formulas in ACQ and ACO. The first motivation is related to Hypercomplex Analysis and the notion of slice regular functions already mentioned in the Introduction. The second motivation to study the fragment of ordered $L$-formulas in ACQ and ACO concerns our Corollary 2.2 and Lemma 2.7. Indeed, let $R\models\textnormal{RCF}$ and consider the extended languages $L_{\mathbb{H}_{R}}$ and $L_{\mathbb{O}_{R}}$, respectively. Then, Corollary 2.2 tells us that $L_{\mathbb{H}_{R}}$-definable and $L_{\mathbb{O}_{R}}$-definable sets are exactly semialgebrac subsets of $\mathbb{H}_{R}^{n}$ and $\mathbb{O}_{R}^{n}$, respectively. On the other hand, by Lemma 2.7, every real algebraic subset of $\mathbb{H}_{R}^{n}$ or $\mathbb{O}_{R}^{n}$ can be described as the solution set of a, in general non-ordered, polynomial with coefficients in $\mathbb{H}_{R}^{n}$ or $\mathbb{O}_{R}^{n}$, respectively. Thus, the classes of algebraic sets defined by ordered polynomials and of definable sets by means of ordered polynomials in $\mathbb{H}_{R}^{n}$ or $\mathbb{O}_{R}^{n}$ constitute subclasses of those sets studied in real algebraic geometry, thus the developement of proper quaternionic and octonionic techniques are of special interest.
Definition 3.3.
Let $X\subset\mathbb{H}_{R}^{n}$ or $X\subset\mathbb{H}_{R}^{n}$. We say that $X$ is an algebraic subset of $\mathbb{H}_{R}^{n}$ or of $\mathbb{O}_{R}^{n}$, respectively, if $X$ can be described as the common solution set of a finite number of ordered polynomial equations with coefficients in $\mathbb{H}_{R}^{n}$ or $\mathbb{O}_{R}^{n}$, respectively. In other words, $X$ is an algebraic subset of $\mathbb{H}_{R}^{n}$ or of $\mathbb{O}_{R}^{n}$, respectively, if
$X$ :=$:=\{(q_{1},\dots,q_{n})\in\mathbb{H}_{R}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(q_{1},\dots,q_{n})=0\}\quad or$
$X$ :=$:=\{(o_{1},\dots,o_{n})\in\mathbb{O}_{R}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(o_{1},\dots,o_{n})=0\},$
respectively, for some ordered polynomials $p_{s}(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ with coefficients in $\mathbb{H}_{R}$ or in $\mathbb{O}_{R}$, respectively, for every $s\in\{1,\dots,\ell\}$.
As a consequence of model completeness for ACQ and ACO next definition is well defined.
Definition 3.4.
Let $R_{1},R_{2}\models\textnormal{RCF}$ such that $R_{1}\preceq R_{2}$. Denote by $q:=(q_{1},\dots,q_{n})$ and $o:=(o_{1},\dots,o_{n})$. Let $X=\{q\in\mathbb{H}_{R_{1}}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(q)=0\}$ or $X=\{o\in\mathbb{O}_{R_{1}}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(o)=0\}$ be an algebraic subset of $\mathbb{H}_{R_{1}}^{n}$ or $\mathbb{O}_{R_{1}}^{n}$, respectively. Denote by $X_{\mathbb{H}_{R_{2}}}:=\{q\in\mathbb{H}_{R_{2}}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(q)=0\}\subset\mathbb{H}_{R_{2}}^{n}$ the extension of $X$ to $\mathbb{H}_{R_{2}}$ or $X_{\mathbb{O}_{R_{2}}}=\{o\in\mathbb{O}_{R_{2}}^{n}\,|\,\bigwedge_{s=1}^{\ell}p_{s}(o)=0\}\subset\mathbb{O}_{R_{2}}^{n}$ the extension of $X$ to $\mathbb{O}_{R_{2}}$, respectively.
Let $R\models\textnormal{RCF}$. Let $p(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ be an ordered polynomial with coefficients in $\mathbb{H}_{R}$ and $p^{\prime}(\mathtt{o}_{1},\dots,\mathtt{o}_{n},)$ be an ordered polynomial with coefficients in $\mathbb{O}_{R}$. Denote by $\mathcal{Z}_{\mathbb{H}_{R}}(p):=\{(q_{1},\dots,q_{n})\in\mathbb{H}_{R}^{n}\,|\,p(q_{1},\dots,q_{n})=0\}$ and $\mathcal{Z}_{\mathbb{O}_{R}}(p^{\prime}):=\{(o_{1},\dots,o_{n})\in\mathbb{O}_{R}^{n}\,|\,p^{\prime}(o_{1},\dots,o_{n})=0\}$ be the quaternion and octonion zero loci of $p$ and $p^{\prime}$, respectively.
Remark 3.5.
Observe that, by Lemma 2.1, after considering real coordinates $(\mathtt{x}_{s,0},\dots,\mathtt{x}_{s,3})$ in the quaternion case and real coordinates $(\mathtt{x}_{s,0},\dots,\mathtt{x}_{s,7})$ in the octonion case such that $\mathtt{q}_{s}=\mathtt{x}_{s,0}+i\mathtt{x}_{s,1}+j\mathtt{x}_{s,2}+k\mathtt{x}_{s,3}$ and $\mathtt{o}_{s}=\mathtt{x}_{s,0}+\sum_{t=1}^{7}e_{s,t}\mathtt{x}_{s,t}$ for every $s\in\{1,\dots,n\}$ we have that both $\mathcal{Z}_{\mathbb{H}_{R}}(p)$ and $\mathcal{Z}_{\mathbb{O}_{R}}(p^{\prime})$ are real algebraic sets.
Let us generalize properties of ordered polynomials with coefficients in $\mathbb{H}_{R}$ or in $\mathbb{O}_{R}$ originally proved in [GS08, GP11, GP22] in the case $R={\mathbb{R}}$.
Let us introduce a first-order characterization of dimension for algebraic subsets $X$ of $R^{n}$. Let $X\subset R^{n}$ be an algebraic set. Denote by $\mathcal{I}(X)\subset R[\mathtt{x}_{1},\dots,\mathtt{x_{n}}]$ the ideal of polynomials vanishing over $X$. Recall that, by definition, $\dim_{R}(X)$ is the Krull dimension of the ring $R[\mathtt{x}_{1},\dots,\mathtt{x_{n}}]/\mathcal{I}(X)$.
Lemma 3.6.
Fix $L_{R}:=L\cup\{c_{i}\}_{i\in R}$ to be the extension of $L$ with constant symbols coming from $R$. Let $X\subset R^{n}$ be an algebraic set, then $\dim_{R}(X)$ is $L_{R}$-definable.
Proof.
Let $X_{1},\dots,X_{\ell}$ be the irreducible components of $X$. By [BCR98, Theorem 2.8.3(i)], we have $\dim_{R}(X)=\max(\dim_{R}(X_{1}),\dots,\dim(X_{\ell}))$. Observe that, for every $a,b\in R$ with $a\neq b$, $a<b$ if and only if $\exists c(b-a=c^{2})$, thus it is possible to describe the maximum of a finite set without using the symbol $<$. Thus, we are only left to prove that $\dim_{R}(X)$ is $L$-definable for an irreducible algebraic set $X\subset R^{n}$. Let $X\subset R^{n}$ be an irreducible algebraic set and let $\mathcal{I}(X)=(p_{1},\dots,p_{\ell})$. Denote by $\textnormal{Reg}(X):=\{x\in X\,|\,\textnormal{rk}([\frac{\partial p_{s}}{\partial\mathtt{x}_{t}}(x)]_{s=1,\dots,\ell,\,t=1,\dots,n})=n-\dim_{R}(X)\}$ the set of nonsingular points of $X$ (see [BCR98, Definition 3.3.4]). Observe that $\textnormal{Reg}(X)\subset X$ is a non-empty Zariski open set, hence Zariski dense in $X$. In addition, [BCR98, Proposition 3.3.2] shows that we have an equivalent definition of $\dim_{R}(X)$ in terms of nonsingular points, that is:
$\dim_{R}(X)=\max\Big{(}\Big{\{}d\in\{1,\dots,n\}\,\Big{|}\,\exists x\in X\,\Big{(}\textnormal{rk}\Big{(}\Big{[}\frac{\partial p_{s}}{\partial\mathtt{x}_{t}}(x)\Big{]}_{s=1,\dots,\ell,\,t=1,\dots,n}\Big{)}=n-d\Big{)}\Big{\}}\Big{)},$
where $\mathtt{x}:=(\mathtt{x}_{1},\dots,\mathtt{x}_{n})$. Observe that latter equivalent definition of $\dim_{R}(X)$, for an irreducible algebraic set $X\subset R^{n}$, is $L_{R}$-definable since the rank of a matrix can be described in terms of vanishing minors, as desired. ∎
Remark 3.7.
Assume $X\subset R^{n}$ is a ${\mathbb{Q}}$-algebraic set, that is, $X$ is an algebraic that can be described by polynomial equations without coefficients. Up to consider the ${\mathbb{Q}}$-irreducible components of $X$ (see [FG, Definition 2.1.5]) we can supopse $X$ is ${\mathbb{Q}}$-irreducible as in Lemma 3.6. Then, the thesis of Lemma 3.6 can be refined as $\dim_{R}(X)$ is $L$-definable even though, in general,
$\mathcal{I}(X)=\mathcal{I}_{\overline{{\mathbb{Q}}}^{r}}(X)R[\mathtt{x}]\supsetneq\mathcal{I}_{{\mathbb{Q}}}(X)R[\mathtt{x}],$
where $\mathcal{I}_{K}(X):=\mathcal{I}(X)\cap K[\mathtt{x}]$ denotes the ideal of polynomials with coefficients in $K$ vanishing over $X$, for $K={\mathbb{Q}},\overline{{\mathbb{Q}}}^{r}$ (see [FG, Theorem 3.1.2]). Indeed, let $\mathcal{I}_{\mathbb{Q}}(X)=(q_{1},\dots,q_{\ell})$, for some $q_{1}(\mathtt{x}),\dots,q_{\ell}(\mathtt{x})\in{\mathbb{Q}}[\mathtt{x}]$. As above, we may reduce to the case $X\subset R^{n}$ is ${\mathbb{Q}}$-irreducible, that is, $\mathcal{I}_{\mathbb{Q}}(X)$ is a prime ideal of ${\mathbb{Q}}[\mathtt{x}]$. In [FG, Definition 5.1.1], Fernando and Ghiloni introduce the notion of $R|{\mathbb{Q}}$-regular points and [FG, Theorem 5.4.1] they prove
$\textnormal{Reg}^{*}(X):=\Big{\{}x\in X\,\Big{|}\,\textnormal{rk}\Big{(}\Big{[}\frac{\partial q_{s}}{\partial\mathtt{x}_{t}}(x)\Big{]}_{s=1,\dots,\ell,\,t=1,\dots,n}\Big{)}=n-\dim_{R}(X)\Big{\}}\subset\textnormal{Reg}(X)$
is a non empty ${\mathbb{Q}}$-Zarski open subset of $X$. Hence, if $X\subset R^{n}$ is a ${\mathbb{Q}}$-algebraic set, then
$\dim_{R}(X)=\max\Big{(}\Big{\{}d\in\{1,\dots,n\}\,\Big{|}\,\forall\mathtt{x}\,\Big{(}\textnormal{rk}\Big{(}\Big{[}\frac{\partial q_{s}}{\partial\mathtt{x}_{t}}(\mathtt{x})\Big{]}_{s=1,\dots,\ell,\,t=1,\dots,n}\Big{)}\leq n-d\Big{)}\Big{\}}\Big{)},$
where $\mathtt{x}:=(\mathtt{x}_{1},\dots,\mathtt{x}_{n})$. Since vanishing conditions on minors are equations over ${\mathbb{Q}}$, we deduce that $\dim_{R}(X)$ is $L$-definable, as desired. For more details about real ${\mathbb{Q}}$-algebraic geometry we refer to [FG, GS23].
Corollary 3.8.
Let $R_{1},R_{2}\models\textnormal{RCF}$ such that $R_{1}\preceq R_{2}$. Let $X\subset\mathbb{H}_{R_{1}}^{n}$ or $X\subset\mathbb{O}_{R_{1}}^{n}$ be an algebraic set. Then, $\dim_{R_{1}}(X)=\dim_{R_{2}}(X_{R_{2}})$.
Proof.
Let $L_{\mathbb{H}_{R_{1}}}:=L\cup\{c\}_{c\in\mathbb{H}_{R_{1}}}$ and $L_{\mathbb{O}_{R_{1}}}:=L\cup\{c\}_{c\in\mathbb{O}_{R_{1}}}$ denote the language $L$ extended with coefficients in $\mathbb{H}_{R_{1}}$ or in $\mathbb{O}_{R_{1}}$, respectively. By Lemma 3.6, after identifying $X$ as a real subset of $R_{1}^{4_{n}}$ or $R_{1}^{8n}$, respectively, then $\dim_{R_{1}}(X)=d$ can be expressed as a first-order formula in the language $L_{R_{1}}$. By Lemma 2.7 and Remark 2.9, we obtain that $\dim_{R_{1}}(X)=d$ can be expressed as a first-order $L_{\mathbb{H}_{R_{1}}}$-formula or as a first-order $L_{\mathbb{O}_{R_{1}}}$-formula, respectively. Then, by model completeness of ACQ and ACO, namely by Theorem 2.4, we deduce that $\dim_{R_{1}}(X)=\dim_{R_{2}}(X_{R_{2}})$, as desired. ∎
3.2. Properties of ordered polynomials
Let $R\models\textnormal{RCF}$. Let us deduce some properties on the real dimension of the zero locus of ordered polynomials.
Theorem 3.9 (Fundamental Theorem of Algebra).
Let $p(\mathtt{q})$ and $p^{\prime}(\mathtt{o})$ be non-constant ordered polynomials with coefficients in $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$, respectively. Then, $\mathcal{Z}_{\mathbb{H}_{R}}(p)\subset\mathbb{H}_{R}$ is a non-empty finite union of isolated points and non-intersecting $2$-spheres and $\mathcal{Z}_{\mathbb{O}_{R}}(p^{\prime})\subset\mathbb{O}_{R}$ is a finite union of isolated points and non-intersecting $6$-spheres.
Proof.
Let us prove that $\mathcal{Z}(p)\subset\mathbb{H}_{R}$ is a non-empty finite union of isolated points and disjoint spheres $\mathbb{S}^{2}\subset\mathbb{H}_{R}$, the proof for $\mathcal{Z}(p^{\prime})$ is similar. Observe that $\mathcal{Z}(p)\neq\emptyset$ because the statement “Every polynomial of positive degree $\leq d$ has a root” can be expressed as a first-order $L$-sentence $\rho_{d}$ for every $d\in{\mathbb{N}}$. Since $\mathbb{H}_{{\mathbb{R}}}\models\rho_{d}$ for every $d\in{\mathbb{N}}$ by [GSV08, Theorem 1.3], model completeness of ACQ, namely Theorem 2.5, ensures that $\mathbb{H}_{R}\models\rho_{d}$ for every $d\in{\mathbb{N}}$ as well. Let $\beta_{d}(\mathtt{q})$ and $\sigma_{d}(\mathtt{q})$ denote the following first-order $L_{\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}}$-formulas:
$\beta_{d}(\mathtt{q},a)$ $:\,\exists r\in R_{+}\,\forall q\Big{(}\Big{(}(q\in B_{R}(\mathtt{q},r))\vee q\neq\mathtt{q}\Big{)}\rightarrow\sum_{i=0}^{d}q^{i}a_{i}\neq 0\Big{)},$
$\sigma_{d}(\mathtt{q},a)$ $:\,\exists c\in R,\,\exists r\in R_{+}\,\Big{(}\mathtt{q}\in S_{R}(c,r)\wedge\forall q\Big{(}(q\in S_{R}(c,r))\rightarrow\sum_{i=0}^{d}q^{i}a_{i}=0\Big{)}\Big{)},$
$\tau_{d}(a)$ $:\,\forall c_{1},c_{2}\in R\,\,\forall r_{1},r_{2}\in R_{+}\,\Big{(}\Big{(}\Big{(}\forall q((q\in S_{R}(c_{1},r_{1}))\rightarrow\sum_{i=0}^{d}q^{i}a_{i}=0)\Big{)}\vee$
$\Big{(}\forall q((q\in S_{R}(c_{2},r_{2}))\rightarrow\sum_{i=0}^{d}q^{i}a_{i}=0)\Big{)}\Big{)}\rightarrow S_{R}(c_{1},r_{1})\cap S_{R}(c_{2},r_{2})=\emptyset\Big{)},$
where $a:=(a_{1},\dots,a_{d})\in\mathbb{H}_{R}^{d}$, $R_{+}:=\{q\in\mathbb{H}_{R}\,|\,q\in R,q\neq 0,\exists p(q=p^{2})\}$, $B_{R}(\mathtt{q},r)$ denotes the ball in $\mathbb{H}_{R}$ of center $\mathtt{q}$ and radius $r\in R_{+}$ and $S_{R}(c,r)$ the sphere in $\mathbb{H}_{R}$ of center $c$ and radius $r\in R_{+}$. Define the $L_{\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}}$-sentence $\phi_{d}$ as follows:
$\phi_{d}:\,\forall a_{0}\dots\forall a_{d}\forall\mathtt{q}\Big{(}\big{(}\sum_{i=0}^{d}\mathtt{q}^{i}a_{i}=0\big{)}\rightarrow\big{(}\beta_{d}(\mathtt{q})\vee(\sigma_{d}(\mathtt{q})\wedge\tau_{d})\big{)}\Big{)}.$
Observe that $\phi_{d}$ is the sentence expressing the following statement: “The zero locus of each ordered polynomial of degree $\leq d$ consists of a finite number of isolated points and disjoint spheres centered at real values.”. Observe that the number of points and spheres is finite because the zero locus $\mathcal{Z}(p)\subset\mathbb{H}_{R}$ of $p$ is also an algebraic set in real coordinates, so it has a finite number of real irreducible components, see [BCR98, Theorem 2.8.3]. Observe that $\mathbb{H}\models\phi_{d}$ for every $d\in{\mathbb{N}}$ by [GSV08, Theorem 1.3] and $\phi_{d}$ is a $L_{\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}}$-sentence by Remark 3.7 and Lemma 2.7. Thus, completeness of ACQ, namely Theorem 2.5, ensures that $\mathbb{H}_{R}\models\phi_{d}$ for every $d\in{\mathbb{N}}$ and for every $R\models\textnormal{RCF}$, as desired. ∎
Remark 3.10.
There are non-ordered polynomials with coefficients in $\mathbb{H}_{R}$ and $\mathbb{O}_{R}$ having empty zero locus. Consider the polynomial $p(q):=\frac{1}{16}(q-iqi-jqj-kqk)^{2}+1$. In real coordinates of $\mathbb{H}_{R}$ the polynomial $p(q)$ corresponds to $x_{0}^{2}+1$, which has no real roots in $R^{4}$.
Remark 3.11.
The class of algebraic sets as in Definitions 3.3&3.4 is strictly smaller than the class of sets that can be defined by finite systems of non-ordered polynomial equations. Consider $\mathbb{H}_{R}$, a similar argument works for $\mathbb{O}_{R}$. Consider the set
$X:=\Big{\{}q=x_{0}+ix_{1}+jx_{2}+kx_{3}\in\mathbb{H}_{R}\,\Big{|}\,x_{0}=\frac{1}{4}\Big{(}q-iqi-jqj-kqk\Big{)}=0\Big{\}}$
defined by a non-ordered polynomial equation. Observe that its real dimension is $3$ and it is not compact in the euclidean topology of $\mathbb{H}_{R}$. However, as a consequence of Theorem 3.9, algebraic subsets of $\mathbb{H}_{R}$ have real dimension $\leq 2$ and they are compact with respect to the euclidean topology of $\mathbb{H}_{R}$.
Here we propose an extension of [GP22, Propositions 3.15 & 3.17] for every model of ACQ and ACO. We point out that our proofs are direct consequences of completeness and Corollary 3.8, although the original proofs over $\mathbb{H}$ and $\mathbb{O}$ naturally extend over quaternions and octonions over any real closed field.
Proposition 3.12.
Let $R\models\textnormal{RCF}$. Let $p(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$ and $p^{\prime}(\mathtt{o}_{1},\dots,\mathtt{o}_{n})$ be non-constant ordered polynomials with coefficients in $\mathbb{H}_{R}$ and in $\mathbb{O}_{R}$, respectively. Then
$24(n-1)\leq\dim_{R}(\mathcal{Z}_{\mathbb{H}_{R}}(p))\leq 4n-$ 2
and
$28(n-1)\leq\dim_{R}(\mathcal{Z}_{\mathbb{O}_{R}}(p^{\prime}))\leq 8n-$ 2 .
Proof.
We prove the quaternionic case, the same strategy works for octonions as well. By Lemma 3.6 and Lemma 2.7, the dimension of an algebaic subset of $\mathbb{H}_{R}^{n}$ is $L_{\mathbb{H}_{R}}$-definable and, if the starting algebraic ia the common solution set of $L$-polynomials, then its dimension is $L_{\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}}$-definable by Remark 2.9. Let $p(\mathtt{q}):=\sum_{|\alpha|\leq d}\mathtt{q}_{1}^{\alpha_{1}}\dots\mathtt{q}_{n}^{\alpha_{n}}a_{\alpha}$ where $\mathtt{q}:=(\mathtt{q}_{1},\dots,\mathtt{q}_{n})$, $\alpha\in{\mathbb{N}}^{n}$ and $a_{\alpha}\in\mathbb{H}_{R}$. Let $\psi_{k}(a)$ be the first order $L_{\mathbb{H}_{R}}$-formula, depending on the string of coefficients $a$ of $p(\mathtt{q})$, stating: “The zero locus of the ordered polynomial $p(\mathtt{q})$ has real dimension $k$.”. Let $\Lambda:=\{m\in{\mathbb{N}}^{n}\,|\,|m|\leq d\}$ and let $a:=(a_{\alpha})_{\alpha\in\Lambda}$ be a string of elements of $\mathbb{H}_{R}$. Consider the $L$-sentence $\phi_{d}$ defined as:
$\phi_{d}:\quad\forall a\Big{(}\psi_{4(n-1)}(a)\vee\psi_{4n-3}(a)\vee\psi_{4n-2}(a)\Big{)}.$
Observe that $\phi_{d}$ is a $L_{\mathbb{H}_{\overline{{\mathbb{Q}}}^{r}}}$-sentence stating: “Every ordered polynomial $p(\mathtt{q})$ of degree $\leq d$ satisfies $24(n-1)\leq\dim_{R}(\mathcal{Z}_{\mathbb{H}_{R}}(p))\leq 4n-$ 2.”. By [GP22, Propositions 3.15 & 3.17] we have $\mathbb{H}\models\phi_{d}$ for every $d\in{\mathbb{N}}$, thus completeness of ACQ, namely Theorem 2.5, ensures that $\mathbb{H}_{R}\models\phi_{d}$ for every $d\in{\mathbb{N}}$ and for every $R\models\textnormal{RCF}$, as desired. ∎
Remark 3.13.
[…]
3.3. Failure of Quantifier Elimination for the fragment of ordered formulas
This last section is devoted to study the fragment of ACQ and ACO consting of those $L$-formulas that can be written only by means of ordered $L$-terms. By ordered $L$-term we mean a first-order $L$-term in which only ordered $L$-polynomials are involved. We refer to the fragment of ordered $L$-formulas as the fragments of ACQ and ACO in which only ordered $L$-terms occur. In principle, since the class of ordered polynomials have much strict algebraic properties (see Theorem 3.9 and Proposition 3.12) it is worth to study properties of this subclass of $L$-formulas. Unfortunately, quantifier elimination does not hold for the fragment of ordered $L$-formulas as well since every $L$-formula $\varphi$ is equivalent, modulo ACQ and ACO, to an ordered $L$-formula $\psi$.
Theorem 3.14.
Every $L$-formula is equivalent modulo ACQ and ACO to an ordered $L$-formula. Equivalently, every $L$-definable subset of $\mathbb{H}_{R}^{n}$ and $\mathbb{O}_{R}^{n}$ is ordered $L$-definable.
Proof.
By induction on first-order $L$-formulas it suffices to prove the statement for atomic $L$-formulas. We only prove the quaternionic case, the same procedure works for octonions as well. Let $p(\mathtt{q})$ be a $L$-polynomial and consider the atomic formula $p(\mathtt{q})=0$. Let us prove by induction on the number $a$ of monomials defining $p(\mathtt{q})$ that are not of the form $\mathtt{q}_{1}^{\alpha_{1}}\cdot\mathtt{q}_{2}^{\alpha_{2}}\cdot\,\dots\,\cdot\mathtt{q}_{n}^{\alpha_{n}}$, for some $\alpha:=(\alpha_{1},\dots,\alpha_{n})\in{\mathbb{N}}^{n}$. If $a=0$, then $p(\mathtt{q})$ is an ordered $L$-polynomial and there is nothing to prove. Suppose $p(\mathtt{q})$ has $a>0$ monomials that are not of the form $\mathtt{q}_{1}^{\alpha_{1}}\cdot\mathtt{q}_{2}^{\alpha_{2}}\cdot\,\dots\,\cdot\mathtt{q}_{n}^{\alpha_{n}}$, for some $\alpha:=(\alpha_{1},\dots,\alpha_{n})\in{\mathbb{N}}^{n}$. Let
$\mathtt{q}_{b_{1}}^{\beta_{b_{1}}}\cdot\mathtt{q}_{b_{2}}^{\beta_{b_{2}}}\cdot\,\dots\,\cdot\mathtt{q}_{b_{\ell}}^{\beta_{b_{\ell}}}$
be one of such monomials, for some $\ell\in{\mathbb{N}}^{*}$, $\beta_{1},\beta_{2},\dots,\beta_{\ell}\in{\mathbb{N}}$ and $b_{1},b_{2}\dots,b_{\ell}\in\{1,\dots,n\}$ with $b_{h}\neq b_{h+1}$ for every $h\in\{1,\dots,\ell-1\}$. Let $h^{\prime}\in\{1,\dots,\ell\}$ such that $b_{h}<b_{h+1}$ for every $h\in\{1,\dots,h^{\prime}-1\}$ and $b_{h^{\prime}}>b_{h^{\prime}+1}$. Consider the monomial
$\mathtt{q}_{b_{1}}^{\beta_{b_{1}}}\cdot\mathtt{q}_{b_{2}}^{\beta_{b_{2}}}\cdot\,\dots\,\cdot\mathtt{q}_{b_{h}}^{\beta_{b_{h}}}\cdot\mathtt{r}_{1}^{\beta_{b_{h+1}}}\cdot\mathtt{r}_{2}^{\beta_{b_{h+2}}}\cdot\,\dots\,\cdot\mathtt{r}_{\ell-h^{\prime}}^{\beta_{b_{\ell}}}$
in ordered variables $\mathtt{q}_{1},\dots,\mathtt{q_{n}},\mathtt{r}_{1},\dots,\mathtt{r}_{\ell-h}$ and define $p^{\prime}(\mathtt{q}_{1},\dots,\mathtt{q_{n}},\mathtt{r}_{1},\dots,\mathtt{r}_{\ell-h^{\prime}})$ as the polynomial obtained by substituting the monomial $\mathtt{q}_{b_{1}}^{\beta_{b_{1}}}\cdot\mathtt{q}_{b_{2}}^{\beta_{b_{2}}}\cdot\,\dots\,\cdot\mathtt{q}_{b_{\ell}}^{\beta_{b_{\ell}}}$ in $p(\mathtt{q})$ with the above ordered monomial. Then, $p^{\prime}(\mathtt{q}_{1},\dots,\mathtt{q_{n}},\mathtt{r}_{1},\dots,\mathtt{r}_{\ell-h})$ has $a-1$ monomials that are not of the form
$\mathtt{q}_{1}^{\alpha_{1}}\cdot\mathtt{q}_{2}^{\alpha_{2}}\cdot\,\dots\,\cdot\mathtt{q}_{n}^{\alpha_{n}}\cdot\mathtt{r}_{1}^{\alpha_{n+1}}\cdot\dots\cdot\mathtt{r}_{n+\ell-h^{\prime}}^{\alpha_{\ell-h^{\prime}}},$
for some $\alpha^{\prime}:=(\alpha_{1},\dots,\alpha_{n},\alpha_{n+1},\dots,\alpha_{n+\ell-h^{\prime}})\in{\mathbb{N}}^{n+\ell-h^{\prime}}$.
Consider the ordered $L$-formula $\phi(\mathtt{q},\mathtt{r})$ defined as follows:
$(p^{\prime}(\mathtt{q},\mathtt{r})=0)\wedge\Big{(}\bigwedge_{h^{\prime}=1}^{\ell-h}(\mathtt{r}_{h^{\prime}}-\mathtt{q}_{b_{h+h^{\prime}}}=0)\Big{)}.$
Observe that the formulas $p(\mathtt{q})=0$ and $\exists r\phi(\mathtt{q},r)$ are elementary equivalent modulo ACQ, as desired. ∎
Remark 3.15.
A symilar statement of Theorem 3.14 holds for $L^{\prime}$-formulas, where $L^{\prime}$ denotes $L\cup\{q_{i}\}_{i\in\mathbb{H}_{R}}$ and $L\cup\{o_{i}\}_{i\in\mathbb{O}_{R}}$, respectively. Indeed, it suffices to apply the same substitution procedure to the coefficients that may occur in non-ordered $L^{\prime}$-monomials between two variables.
As a direct consequence of Theorem 3.14 we deduce the following result.
Corollary 3.16.
The $L$-theories of ACQ and ACO do not admit quantifier elimination for the fragment of ordered $L$-formulas.
Latter result shows that, restricting our interest on ordered $L$-formulas is not useful to understand which $L$-formulas admit quantifier elimination. Let $L^{\prime}$ denotes either the lenguage $L_{\mathbb{H}_{R}}$ or $L_{\mathbb{O}_{R}}$. We observe that extending the language to $L^{\prime}$ the set of ordered quantifier free $L^{\prime}$-definable subsets of $\mathbb{H}_{R}^{n}$ or $\mathbb{O}_{R}^{n}$, respectively, is strictly contained in the set of quantifier free $L_{\mathbb{H}_{R}}$-definable subsets of $\mathbb{H}_{R}^{n}$, as it is shown by the following example.
Example 3.17.
Let $R\models RCF$. Consider the subset $\mathcal{Z}_{\mathbb{H}_{R}}(\mathtt{p})\subset\mathbb{H}_{R}$, with $p:=\mathtt{q}-i\mathtt{q}i-j\mathtt{q}j-k\mathtt{q}k$, as in Remark 3.11. By Theorem 3.9, the zero set of an ordered polynomial in one variable has real dimension 00 or $2$ as real agebraic subsets of ${\mathbb{R}}^{4}$. As a consequence, those subsets of $\mathbb{H}_{R}$ defined by quantifier free $L$-formulas have real dimension $d\in\{0,1,2,4\}$ as locally closed subsets of $R^{4}$. Since $\dim_{R}(X)=3$, we deduce that $X$ is quantifier free definable but it is not ordered quantifier free definable.